OpenStudy (anonymous):

Write an equation in slope-intercept form of the line satisfying the given conditions. Parallel to the line x +5y=6.containing the point(0,0)

OpenStudy (***[isuru]***):

when 2 lines r parallel .. the slopes r equal and the standered form of a straight line is y = mx + b where m = slope b = intercept.. since the line is parallel to x +5y=6 which is 5y = 6 - x y = (-1/5) x + 6/5 the slope of the line u want should be the same.. so u can write the equation of the line u want as.. y = (-1/5) x + b now to find b u can use the given coordinates for that ( 0 , 0 ) plug them in the equation u've got y = (-1/5)x + b 0 (-1/5)0 + b b= 0 therefor the line parallel to x +5y=6 and goes through (0,0) is y = (-1/5)x got it :) ?

OpenStudy (anonymous):

I am grateful.thanks